The management at MooseWoods Park is trying to decide whether to replace their portable toilets with flush toilets at the very popular MooseWoods Campground, as well as how many flush toilets to install, if they do choose to install them. They concluded that they will install flush toilets if the total surplus is greater than $50

in relation to the number of flush toilets at the equilibrium and the cost per campsite per night. Installing flush toilets will involve an increase in the cost per campsite per night.

Researchers have calculated the demand and supply curves for flush toilets at this park and discovered that they are both perfectly straight lines. At $35

per campsite per night, no one will want flush toilets. At $0

per campsite per night, campers will demand seven flush toilets.

Managers are willing to supply no flush toilets at the current campsite cost of $5

per night, but are willing to supply eight flush toilets at $25

per campsite per night.

Please use the supply, demand and total surplus information to answer the three questions.

What is the number of flush toilets desired at the equilibrium point?

toilets

What is the total surplus?

$

Will the managers install the number of flush toilets determined in the first question?

not enough information to determine

yes

no
Answer #1

A. Number of flush toilets =4

Draw the demand and supply curves from the given information.

301 Price 334 30 27 24 Supply 21- -18 -15 Demand Quantity

B. Total surplus is the area of the triangle ABC

Total surplus= 1/2*(35-5)*(4)= 1/2*30*4= $60

C. Yes

Reason-Since the total surplus $60 is greater than $50, managers will install the flush toilets.